OpenStudy (anonymous):

What is the mean of the following set of numbers? 9 6/8 , 3.9 , 8.4 A. 3.9 B. 7.35 C. 2.8 D. 11.025

sam (.sam.):

B. 7.35

OpenStudy (agreene):

B 7.35, Remeber Arithmetic Mean is the sum of the numbers of the group divided by the number of members of the group.

OpenStudy (agreene):

Remember *

OpenStudy (anonymous):

B. 7.35

sam (.sam.):

\[\frac{\frac{39}{4}+3.9+8.4}{3}=7.35\]